OpenStudy (anonymous):

If the sin 90° = 1, then the cos 0° = _____.

OpenStudy (anonymous):

I know sin(Ø-90)=cos(Ø) But I was wondering if the 1 is carried or changed.

OpenStudy (inkyvoyd):

is that the only identity you are given?

OpenStudy (anonymous):

Im only given this: If the sin 90° = 1, then the cos 0° = _____. And a few answers: 0 because the angles are complementary 1 over 22 because the angles are complementary square root 3 over 2 because the angles are complementary 1 because the angles are complementary

OpenStudy (inkyvoyd):

do you know what complementary angles are?

OpenStudy (anonymous):

Two angles where both or either equal 90°?

OpenStudy (inkyvoyd):

no, it's two angles that add up to 90 degrees. the identity you gave was actually incorrect. you switched the terms inside the sine function, it should be sin(90-Ø)=cos(Ø)

OpenStudy (anonymous):

ok.

OpenStudy (inkyvoyd):

so if sin(90)=1, then what does Ø equal?

OpenStudy (anonymous):

cos(0)=?

OpenStudy (inkyvoyd):

well, Ø would equal 0, but yes, sin(90)=cos(0) but what does sin(90) equal?

OpenStudy (inkyvoyd):

*also equal

OpenStudy (anonymous):

1

OpenStudy (inkyvoyd):

and you have your answer :)

OpenStudy (anonymous):

:) thanks